# Topological Phases of Many-Body Localized Systems: Beyond Eigenstate Order

### v3

@@ Appendix A From ALT phases to QCA
We formalize our definitions of “MBL Hamiltonian” and “short-ranged entanglement” thereof in Appendix A.1. The main point of this section and Appendix A.2 is that these physical notions reproduce the formal requirements of [56, Theorem II.4]. Thus, we realize abstract restrictions on commuting projector models as physical properties of localized systems. [56, Theorem II.4] then states that the LIOMs of the MBL Hamiltonian can be prepared by a QCA. We make an equivalent proof in Appendix A.2. The QCA in question is not unique, and we characterize the redundancy in Appendix A.3. We also discuss localized phases of driven systems, which we also characterize (non-uniquely) by parameterized families of QCA.
@@ A.1 Setup
Here, $B_{r}(S)$ is the ball of radius $r$ around the set $S$. Properly, a QCA should be a projective unitary, but the distinction will not be important here. Reference [51] defines a commuting model to be “without intrinsic topological order” when the LIOMs $\tau^{z}_{j}$ can be prepared from single site operators
@@ A.2 Existence
In fact, after a redefinition of LIOMs to another equivalent set, we will have that $V^{\dagger}\tau^{z}_{j}V$ is a single site operator.
@@
Note that every vertex in $B$ has finite degree by the assumption that $K$ and $J$ are locally finite. Further observe that the graph $B$ splits into distinct connected components labelled by primes $p$ such that $N_{j}=M_{k}=p$. Thus, we can consider each component separately. Henceforth, we assume that we are working in a given component.
@@
is a sum of products of the fixed $\sigma^{z}_{k}$ operators and the micromotion operator $V({\vec{\theta}})$ is a continuous parameterized family of QCA. By a similar proof to the above, we see that this definition is equivalent to the existence of a maximal set of LIOMs $\tau^{z}_{j}({\vec{\theta}})$ with explicit continuous $\vec{\theta}$} dependence and a family of conjugate $\tau^{x}_{j}({\vec{\theta}})$ operators also with continuous $\vec{\theta}$} dependence. Alternatively, the definition is also equivalent to the quasienergy operator [129, 103, 130, 131, 132] having a complete set of LIOMs [104] and conjugate operators which create or destroy excitations locally.
@@ A.3 Non-uniqueness
where $N$ is the Hilbert space dimension. Thus, multiplication by $P_{m}(\theta)$ can take $\widetilde{W}_{1}[V^{\prime}]$ to any allowed value, and we conclude that all loops of zero dimensional QCA (loops of projective unitaries) define trivial localized phases.
@@ Appendix B The $\Omega$-spectrum property
In this appendix, we provide further arguments in support of this conjecture. In Appendix B.1, we recall that if loops of QCA can be put on open boundary conditions such that they act by QCA on their edges—known as blending with the identity—then loops of QCA can all be parameterized as finite depth circuits of lower-dimensional QCA. We also argue that this condition holds for loops of one-dimensional QCA. Then, in Appendix B.2, we construct an explicit deformation retract from the space of circuits to the image of the swindle map. As the swindle map is injective, this shows
@@ B.1 Blending and circuits
where $\mathrm{supp}(A)$ is the smallest subset of ${\mathbb{Z}}^{d}$ on which the operator $A$ acts non-trivially, $x_{l,r}$ are positions along the first axis in ${\mathbb{Z}}^{d}$, and $\alpha(\cdot)$ is the map on local operators defined by conjugation by the locality-preserving unitary defining $\alpha$. That is, $\gamma$ agrees with $\alpha$ to the left of some cut, and agrees with $\beta$ to the right of the cut. We can similarly define a blend $\gamma({\vec{\theta}})$ between parameterized families of QCA $\alpha({\vec{\theta}})$ and $\beta({\vec{\theta}})$. Two (families of) QCA are said to blend into each other if a blend exists between them.
@@
We can split $\gamma$ into a depth two circuit by iteratively splitting off gates of the form $\tilde{\gamma}_{x}\tilde{\gamma}_{x+1}^{-1}$ or $\tilde{\gamma}^{-1}_{x}\tilde{\gamma}_{x+1}$. (The gates are themselves QCA in d1d-1 dimensions.) First, observe that we can split $\gamma$ into a product of two semi-infinite QCA
@@
Nonetheless, we expect that the path-space $P\mathrm{QCA}_{d}$ can be continuously parameterized in terms of a strictly local path of circuits in dimensions higher than one too. The loop space $\Omega\mathrm{QCA}_{d}$ can then be parameterized in terms of circuits which begin and end at the identity. Note that this does not require that the individual gates must return to the identity, only that they cancel at the end of the path.
@@ C.1.1 Unstable invariants
We first recall that connected components of symmetric projective unitaries ($n=0$) are labelled by one-dimensional representations of $G$. Given a symmetric projective unitary $V$, we can arbitrarily choose a proper unitary $\tilde{V}$ in the preimage of $V$ under the quotient by global phases. The fact that $V$ is symmetric only implies that $\tilde{V}$ is symmetric up to a global phase,
@@ C.2 Functorality
What’s more, any symmetric $V\in\mathrm{QCA}_{d}^{\mathcal{R}}$ can also be regarded as being symmetric with respect to the $H$ action $\rho\phi$. Thus, we have a continuous inclusion map
@@ C.3 Application to the QP pump
We can also see the functorality of Appendix C.2 in action. Suppose that we explicitly break the symmetry. Then the charge sector $1\leftrightarrow x$ should be mapped to having charge 0. The new invariant can be found by evaluating the rational function $\widetilde{W}_{1}^{Q}[v]$ at $x=1$. We find
@@
Indeed, we believe the odd $W$ phase without symmetry to be the two-tone-driven phase of Refs. [29, 30]. While the solvable limit of the model in Ref. [29] is highly singular, which makes computing invariants from integral formulae inconvenient, the edge phenomenology described in Ref. [29] is consistent with what we expect from having $\widetilde{W}_{1}\neq 0\bmod 1$. Namely, the edge dynamics was argued to be anomalous in that it could not be achieved by a zero-dimensional localized [in the sense that a continuous micromotion operator exists as in Eq. (133)] system which did not occur at the edge of an extended one-dimensional system. As the model in Ref. [29] is based around a version of the AKLT model, which can be trivialized when stacked with itself, it is thus reasonable to suppose that $\widetilde{W}_{1}=1/2\bmod 1$ for that model, as this is the only element of $\mathbb{Q}$$\mathbb{Z}$} of order 2. Thus, the abstract classification reveals a relationship between the interacting QP pump and Refs. [29, 30] which was not clear previously. It also extends the arguments for stability in Ref. [29] to the nonperturbative regime, where a picture of emergent symmetry (as employed in that reference) becomes inapplicable.

### v4

@@ Appendix A From ALT phases to QCA
We formalize our definitions of “MBL Hamiltonian” and “short-ranged entanglement” thereof in Appendix A.1. The main point of this section and Appendix A.2 is that these physical notions reproduce the formal requirements of [56, Theorem II.4]. Thus, we realize abstract restrictions on commuting projector models as physical properties of localized systems. [56, Theorem II.4] then states that the LIOMs of the MBL Hamiltonian can be prepared by a QCA. We make an equivalent proof in Appendix A.2. The QCA in question is not unique, and we characterize the redundancy in Appendix A.3. We also discuss localized phases of driven systems, which we also characterize (non-uniquely) by parametrized families of QCA.
@@ A.1 Setup
Here, $B_{r}(S)$ is the ball of radius $r$ around the set $S$. Properly, a QCA should be a projective unitary, but the distinction will not be important here. Reference [51] defines a commuting model to be “without intrinsic topological order” when the LIOMs $\tau^{z}_{j}$ can be prepared from single-site operators
@@ A.2 Existence
In fact, after a redefinition of LIOMs to another equivalent set, we will have that $V^{\dagger}\tau^{z}_{j}V$ is a single-site operator.
@@
Note that every vertex in $B$ has finite degree by the assumption that $K$ and $J$ are locally finite. Further observe that the graph $B$ splits into distinct connected components labeled by primes $p$ such that $N_{j}=M_{k}=p$. Thus, we can consider each component separately. Henceforth, we assume that we are working in a given component.
@@
is a sum of products of the fixed $\sigma^{z}_{k}$ operators and the micromotion operator $V({\vec{\theta}})$ is a continuous parametrized family of QCA. By a similar proof to the above, we see that this definition is equivalent to the existence of a maximal set of LIOMs $\tau^{z}_{j}({\vec{\theta}})$ with explicit continuous $\vec{\theta}$} dependence and a family of conjugate $\tau^{x}_{j}({\vec{\theta}})$ operators also with continuous $\vec{\theta}$} dependence. Alternatively, the definition is also equivalent to the quasienergy operator [129, 103, 130, 131, 132] having a complete set of LIOMs [104] and conjugate operators which create or destroy excitations locally.
@@ A.3 Non-uniqueness
where $N$ is the Hilbert space dimension. Thus, multiplication by $P_{m}(\theta)$ can take $\widetilde{W}_{1}[V^{\prime}]$ to any allowed value, and we conclude that all loops of zero-dimensional QCA (loops of projective unitaries) define trivial localized phases.
@@ Appendix B The $\Omega$-spectrum property
In this appendix, we provide further arguments in support of this conjecture. In Appendix B.1, we recall that if loops of QCA can be put on open boundary conditions such that they act by QCA on their edges—known as blending with the identity—then loops of QCA can all be parametrized as finite depth circuits of lower-dimensional QCA. We also argue that this condition holds for loops of one-dimensional QCA. Then, in Appendix B.2, we construct an explicit deformation retract from the space of circuits to the image of the swindle map. As the swindle map is injective, this shows
@@ B.1 Blending and circuits
where $\mathrm{supp}(A)$ is the smallest subset of ${\mathbb{Z}}^{d}$ on which the operator $A$ acts non-trivially, $x_{l,r}$ are positions along the first axis in ${\mathbb{Z}}^{d}$, and $\alpha(\cdot)$ is the map on local operators defined by conjugation by the locality-preserving unitary defining $\alpha$. That is, $\gamma$ agrees with $\alpha$ to the left of some cut, and agrees with $\beta$ to the right of the cut. We can similarly define a blend $\gamma({\vec{\theta}})$ between parametrized families of QCA $\alpha({\vec{\theta}})$ and $\beta({\vec{\theta}})$. Two (families of) QCA are said to blend into each other if a blend exists between them.
@@
We can split $\gamma$ into a depth-two circuit by iteratively splitting off gates of the form $\tilde{\gamma}_{x}\tilde{\gamma}_{x+1}^{-1}$ or $\tilde{\gamma}^{-1}_{x}\tilde{\gamma}_{x+1}$. (The gates are themselves QCA in d1d-1 dimensions.) First, observe that we can split $\gamma$ into a product of two semi-infinite QCA
@@
Nonetheless, we expect that the path-space $P\mathrm{QCA}_{d}$ can be continuously parametrized in terms of a strictly local path of circuits in dimensions higher than one too. The loop space $\Omega\mathrm{QCA}_{d}$ can then be parametrized in terms of circuits which begin and end at the identity. Note that this does not require that the individual gates must return to the identity, only that they cancel at the end of the path.
@@ C.1.1 Unstable invariants
We first recall that connected components of symmetric projective unitaries ($n=0$) are labeled by one-dimensional representations of $G$. Given a symmetric projective unitary $V$, we can arbitrarily choose a proper unitary $\tilde{V}$ in the preimage of $V$ under the quotient by global phases. The fact that $V$ is symmetric only implies that $\tilde{V}$ is symmetric up to a global phase,
@@ C.2 Functoriality
What is more, any symmetric $V\in\mathrm{QCA}_{d}^{\mathcal{R}}$ can also be regarded as being symmetric with respect to the $H$ action $\rho\phi$. Thus, we have a continuous inclusion map
@@ C.3 Application to the QP pump
We can also see the functoriality of Appendix C.2 in action. Suppose that we explicitly break the symmetry. Then the charge sector $1\leftrightarrow x$ should be mapped to having charge 0. The new invariant can be found by evaluating the rational function $\widetilde{W}_{1}^{Q}[v]$ at $x=1$. We find
@@
Indeed, we believe the odd $W$ phase without symmetry to be the two-tone-driven phase of Refs. [29, 30]. While the solvable limit of the model in Ref. [29] is highly singular, which makes computing invariants from integral formulas inconvenient, the edge phenomenology described in Ref. [29] is consistent with what we expect from having $\widetilde{W}_{1}\neq 0\bmod 1$. Namely, the edge dynamics was argued to be anomalous in that it could not be achieved by a zero-dimensional localized [in the sense that a continuous micromotion operator exists as in Eq. (133)] system which did not occur at the edge of an extended one-dimensional system. As the model in Ref. [29] is based around a version of the Affleck-Kennedy-Lieb-Tasaki (AKLT) model, which can be trivialized when stacked with itself, it is thus reasonable to suppose that $\widetilde{W}_{1}=1/2\bmod 1$ for that model, as this is the only element of $\mathbb{Q}$$\mathbb{Z}$} of order 2. Thus, the abstract classification reveals a relationship between the interacting QP pump and Refs. [29, 30] which was not clear previously. It also extends the arguments for stability in Ref. [29] to the nonperturbative regime, where a picture of emergent symmetry (as employed in that reference) becomes inapplicable.